# EventSplat: 3D Gaussian Splatting from Moving Event Cameras for Real-time Rendering

### v1

@@ 1 Introduction
    Use of cubic spline trajectory interpolation for assigning camera poses to events at high rates.
@@ 2 Related Work
2.1 Event Cameras for Efficient Vision Applications
@@ 2.3 Novel View Synthesis via Event Data
Motivated by the unique benefits of event-based cameras, such as high dynamic range, low latency, and high temporal resolution, there has been significant research interest in adopting event data in applications like 3D reconstruction and novel view synthesis in cases where traditional RGB cameras underperform [81, 54, 82, 16]. Recent methods explore using NeRFs with event-vision supervision [59, 31, 25, 42] to reconstruct scenes and render them from novel views. While capable of generating high-quality renderings, NeRF-based methods are generally slow to render due to expensive sampling requirements [6, 17, 57, 62]. On the other hand, 3DGS requires RGB data for supervision, and adapting 3DGS for scenarios where the supervision signal comes from an event-based camera has been challenging [66, 77, 72, 11, 70, 68, 80]. In contrast to previous event-based view synthesis methods, we enable the use of event data to supervise 3D Gaussians, allowing for real-time renderings and state-of-the-art quality.
@@ 4.3 Event-to-Video Guided Initialization
Event cameras capture only the pixels that indicate changes in brightness. This results in sparse information that is often insufficient for effective supervision. Thus, reconstructing 3D information, including background details, solely from event data is exceedingly challenging. To simplify the optimization process we propose leveraging the prior knowledge encoded in an event-to-video model. We use this to generate images from event streams that can subsequently be processed by Structure from Motion (SfM) techniques. Although the generated images often contain significant noise and do not accurately reproduce the underlying RGB image values, they retain substantial texture and background information. This information is sufficient to provide initial positions for the Gaussians and improve the 3DGS optimization process.
@@ 4.4 Event Camera Trajectory Interpolation
The camera poses, $P_{k_{\text{start}}}$ and $P_{k_{\text{end}}}$, used to obtain the log images $\hat{L}_{1}$ and $\hat{L}_{2}$ in Eq. 6, are in most cases not given a priori. This is because event camera poses are typically obtained from a tracking or motion estimation system that operates at a fixed rate, which may not be aligned with the rate of event accumulation. Thus, to obtain an estimate of the complete trajectory of the event camera, we adopt an efficient trajectory interpolation method for camera poses. Our work utilizes Cubic Spline Interpolation and Spherical Cubic Spline Interpolation. These generate a smooth path by interpolating between discrete command points using third-degree polynomials. This interpolation technique closely approximates real-world camera motion by maintaining continuity in both velocity and acceleration.
@@ 5.5 Ablations
To demonstrate the effectiveness of the knowledge-based initialization and the camera trajectory estimation with cubic spline interpolation, we conduct experiments by replacing the component by only adding 1) knowledge-based initialization and 2) adding both knowledge-based initialization and cubic spline interpolation.
Table 3: Ablation results for the knowledge-based initialization and cubic spline interpolation.
@@
prior init 18.75 0.788 0.359
prior init + cubic 18.86 0.792 0.363
@@ 5.6 Limitations
Our method, though effective, has some limitations. It is designed exclusively for novel view synthesis in static scenes and cannot handle scenarios with dynamic moving objects. We consider dynamic event-based view synthesis to be an exciting avenue for future research. Additionally, our approach trains 3D Gaussian Splatting scenes using event-accumulated images between two viewpoints, which represent relative intensity images. Consequently, similar to other event-based view synthesis methods, the model cannot directly estimate absolute intensity images, necessitating a linear transformation using evaluation data as a reference. This linear transformation is required only during inference and does not impact the training process.
@@ B.1 Qualitative Results in Synthetic Scenes
B.2 Qualitative Results in Real-World Scenes
@@
B.3 Qualitative Results of Event-to-Video Guided Initialization
Fig. 8 shows qualitative results for the scene 07_ziggy_and_fuzz_hdr. Due to the effective pre-generation of point clouds for objects such as those in the background, we are able to reconstruct fine details.
Figure 8: The generated image (right) with Event-to-Video Guided Initialization shows better quality, particularly in the background, compared to the generated image (left) with random initialization.

### v2

@@ 1 Introduction
    Use of cubic spline trajectory interpolation to assign camera poses to events at high rates.
@@ 2 Related Work
2.1 Event Cameras for Vision Applications
@@ 2.3 Novel View Synthesis via Event Data
Motivated by the unique benefits of event-based cameras, such as high dynamic range, low latency, and high temporal resolution, there has been significant research interest in adopting event data in applications like 3D reconstruction and novel view synthesis in cases where traditional RGB cameras underperform [81, 54, 82, 16]. Recent methods explore using NeRFs with event-vision supervision [59, 31, 25, 42] to reconstruct scenes and render them from novel views. While capable of generating high-quality renderings, NeRF-based methods are generally slow to render due to expensive sampling requirements [6, 17, 57, 62]. On the other hand, 3DGS requires RGB data for supervision, and adapting 3DGS for scenarios where the supervision signal comes from an event-based camera has been challenging [66, 77, 72, 11, 70, 68, 80]. In contrast to previous event-based view synthesis methods, we enable the use of event data to supervise 3DGS, allowing for real-time renderings and state-of-the-art quality using only event data, supported by event-based initial point clouds and effective camera trajectory interpolation.
@@ 4.3 Event-to-Video Guided Initialization
Event cameras capture only the pixels that indicate changes in brightness. This results in sparse information that is often insufficient for effective supervision. Thus, reconstructing 3D information, including background details, solely from event data is exceedingly challenging. To simplify the optimization process we propose leveraging the prior knowledge encoded in an event-to-video model. We use this pretrained model to generate images from event streams that can subsequently be processed by Structure from Motion (SfM) techniques. Although the generated images often contain significant noise and do not accurately reproduce the underlying RGB image values, they retain substantial texture and background information. This information is sufficient to provide initial positions for the Gaussians, improve the 3DGS optimization process and lead to faster convergence compared to basic initialization with random initial points.
@@ 4.4 Event Camera Trajectory Interpolation
The camera poses, $P_{k_{\text{start}}}$ and $P_{k_{\text{end}}}$, used to obtain the log images $\hat{L}_{1}$ and $\hat{L}_{2}$ in Eq. 6, are in most cases not given a priori. This is because event camera poses are typically obtained from a tracking or motion estimation system that operates at a fixed rate, which may not be aligned with the rate of event accumulation. Thus, to obtain an estimate of the complete trajectory of the event camera, we adopt an efficient trajectory interpolation method for camera poses. Our work utilizes Cubic Spline Interpolation and Spherical Cubic Spline Interpolation. These generate a smooth path by interpolating between discrete command points using third-degree polynomials. This interpolation technique closely approximates real-world camera motion by maintaining non-linear continuity in both velocity and acceleration of camera movement.
@@ 5.5 Ablations
To demonstrate the effectiveness of event-to-video guided initialization and camera trajectory estimation with cubic spline interpolation, we conducted ablation experiments by separately adding 1) event-to-video guided initialization and 2) both the initialization and cubic spline interpolation. As shown in Tab. 3, event-to-video guided initialization outperforms random initialization. Furthermore, Fig. 5 shows that our initialization approach results in better image quality (e.g. backgrounds) compared to random initialization. Adding cubic spline interpolation further enhances performance.
Table 3: Ablation results for the event-to-video guided initialization and cubic spline interpolation.
@@
guided init 18.75 0.788 0.359
guided init + cubic 18.86 0.792 0.363
Figure 5: Qualitative comparisons of the images generated by random initialization and event-to-video guided initialization are shown in 07_ziggy_and_fuzz_hdr scene in EDS dataset.
@@ 5.6 Limitations
Our method, though effective, has some limitations. It is designed exclusively for novel view synthesis in static scenes and cannot handle scenarios with dynamic moving objects, which we consider to be an exciting avenue for future research. Additionally, our approach trains 3D Gaussian Splatting on event-accumulated images between two viewpoints, representing relative intensity images. Consequently, similar to other event-based view synthesis methods, the model cannot directly estimate absolute intensity images and requires a linear transformation with evaluation data as a reference. This linear transformation is required only during inference and does not impact the training.
@@ 6 Conclusion and Future Work
Acknowledgments
The work is in part supported by Sony. We would also like to thank Ziyi Wu, Yash Kant, and Umangi Jain, for valuable discussions and support.
@@ B.1 Qualitative Results in Synthetic Scenes
B.2 Qualitative Results on Different Synthetic Datasets
We evaluated our method on the same scenes used in EV-GS [70] from EventNeRF dataset [59] to compare grayscale results. We computed the mean values across 4 scenes(chair, ficus, hotdog and mic). As shown in Tab. 4, our approach outperforms EV-GS in terms of PSNR and SSIM. Furthermore, Fig. 7 shows qualitative results for 4 synthetic scenes in grayscale, demonstrating that the generated images are reconstructed effectively.
Table 4: Quantitative evaluation of mean values across the 4 synthetic scenes from [70].
Metric PSNR $\uparrow$ SSIM $\uparrow$
Ours EV-GS Ours EV-GS
Mean 29.48 26.6 0.959 0.925
Figure 7: Generated images qualitatively comparing our method with ground truth across 4 synthetic scenes.
B.3 Qualitative Results in Real-World Scenes